A vending machine dispenses hot chocolate or coffee. Service time is 30 seconds per cup and is constant. Customers arrive at a mean rate of 61 per hour, and this rate is Poisson-distributed.

a.

Determine the average number of customers waiting in line. (Round your answer to 2 decimal places.)

  Average number of customer   
b.

Determine the average time customers spend in the system. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Average time minutes  
c.

Determine the average number of customers in the system. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Average number customers  
Answer #1
arrivals/time period = λ= 1.016667 /minute
served/time period= μ= 2.00 /minute

a)

average number of customers in queue Lq = λ2/(μ(μ-λ))= 0.53

b)

average time spend in system W                   = 1/(μ-λ)= 1.02

c)

average number of customers in system L = λ/(μ-λ)= 1.03